Which is equivalent to 1/25^1/3?
a0 5^-6
b) 5^-2/3
c) 5^2/3
d) 5^6 ??
Which is equivalent to 2^x-3/8^x?
a) 2^-3x-3
b) 2^-2x-3
c) 2^2x-3
d) 2^4x-3

Answer by stanbon(75887)

Which is equivalent to 1/25^(1/3)
= [5^(-2)]^(1/3) = 5^(-2/3)
----------------------------
a0 5^-6
b) 5^-2/3
c) 5^2/3
d) 5^6 ??
=====================================
Which is equivalent to 2^x-3/8^x
= (2^(x-3))/[(2^3)^x]
= (2^[(x-3)-3x]
= 2^[-2x-3]
-------------------
a) 2^-3x-3
b) 2^-2x-3
c) 2^2x-3
d) 2^4x-3
==============================================
Cheers,
Stan H.